1. __________What is the relationship for the resistance of nichrome wire relating to its length if {r} represents the resistance of one meter stick resistor and {x} represents the number of meter stick resistors connected end to end with jumper wires.

a. Rtotal=xr  

b. Rtotal=r/x

Explain:

2. __________What is the relationship for the resistance of nichrome wire relating to its cross sectional area if {r} represents the resistance of one meter stick resistor and {x} represents the number of meter stick resistors connected in parallel (side to side in a ladder pattern) with jumper wires.

a. Rtotal=xr  

b. Rtotal=r/x

Explain:
Answer #1

(1)

since resistors connected end to end with jumper wires, the resistor are connected in series so

Rtotal = R+R+R ......... x times = xR

R total = xR

R total = xr ( R= r)

option (a) is correct answer.

(2)

since meter stick resistors connected in parallel, the resistors are connected in parallel

1/R total = 1/R+ 1/R+ 1/R .......... x times = x/R

R total = R/x

R total = r/x

option (b) is correct answer